# Exploratory Study into Relations between Cognitive Distortions and Emotional Appraisals

### v1

@@ 1 Introduction
Emotions are expressed through various modalities, including tone of voice, facial expressions, gestures, and language, particularly in written text. This multifaceted expression of emotions has attracted the interest from NLP and computational researchers in recent years Wang et al. (2022); Plaza-del Arco et al. (2024). While discrete emotional states such as anger, joy, and fear are deemed universal and thus form the basis for automated emotion recognition research, a smaller number of studies have explored dimensional models, representing discrete emotions in continuous spaces Plaza-del Arco et al. (2024). Appraisal theories define emotions as responses that arise from an individual’s evaluation of and event’s significance to their personal goals and well-being, emphasizing that the quality and intensity of emotional responses depend on appraisals, which are the subjective interpretations of the situation Moors et al. (2013).
Negative thoughts are a natural part of human experience; however, they can have a more profound impact on individuals with mental disorders, often becoming entrenched, automatic, and emotionally triggering. Cognitive distortions refer to irrationally exaggerated negative assessments of oneself or situations Beck (1963). Cognitive restructuring, also known as cognitive reframing, is a therapeutic intervention designed to encourage a more positive outlook towards situations by addressing these negative thought patterns Clark (2013). This technique involves replacing negative thoughts with more neutral or hopeful “reframed thoughts”, which provide a softer alternative perspective on the situation.
@@
Our aim in this work is to bring together emotion appraisals and cognitive distortions, to explore the link between these two different but related psychological constructs. We begin by training appraisal prediction models to perform automated appraisal annotations on a dataset annotated with cognitive distortions enabling a combined analysis of both constructs. We analyze the distribution of appraisal values for each distortion-appraisal pair individually, and find statistically significant relations between cognitive distortions and appraisal dimensions, suggesting that different distortion patterns may exhibit distinct appraisal profiles. Finally, when comparing the appraisal profiles between original and reframed texts, we observed a considerable positive shift in several appraisal dimensions, further demonstrating the link between the two constructs and supporting the need for combined study of the two areas.
@@ 2.4 Results
Our model performs on par with the results from Troiano et al. (2023). We use the root-mean-squared error (RMSE) as the evaluation metric and report macro-RMSE averaged over all appraisal dimensions on the test set of 1.36 compared to 1.40 reported in the original paper. Furthermore, Figure 2 compares the performance of the two models for each appraisal dimension, with our model out-performing Troiano et al. (2023) for 13 out of 21 appraisal dimensions. Thus, we deem our trained model good enough to be used for automated appraisal annotation in the remainder of the paper.
Figure 2: Appraisal ratings prediction accuracy in RMSE for both our model and the results reported by Troiano et al. (2023).
@@ 3.1 Thinking trap Dataset
Table 1: Distribution of the cognitive distortion labels of our version of the Thinking Trap dataset.
@@ 3.3 Negative groups
One major consideration is the definition of the negative group within this analysis. The relationship between cognitive distortions and emotional appraisal dimensions, as inferred from the Mann-Whitney test, is strongly influenced by how the positive and negative distributions are defined. Although the definition of the positive group p is fixed, the negative group can have different meanings. Therefore, we consider the following three different definitions of the negative group n:
No distortion: in this case, the negative distribution only contains entries without any distortion, i.e., those belonging to the “no distortion” class. This group represents the appraisal profile of texts without cognitive distortions and acts as a global baseline against which we can compare individual distortion profiles.
Exclusive: here, the negative distribution contains entries that do not belong to the given distortion class but belong to other cognitive distortion classes (excluding “no distortion”). By utilizing this negative group, we can identify differences in appraisal values between various distortion classes. This approach enables us to analyze the appraisal profile of a specific distortion in relation to other distortions, rather than comparing it to a global baseline.
All others: in this configuration we combine both settings by including in the negative group all entries that do not belong to the specified distortion class.
@@ 3.4 Results
Figure 3 plots the discretized significance values for two definitions of the negative groups considered in our study. [^6] Please refer to Appendix A for remaining plots. In these plots, a cell colored white represents a statistically significant difference in appraisal values between the positive and negative groups, while a black cell indicates no statistically significant difference.
@@ 4 Distortions and Corresponding Appraisal Profiles
Figure 5: Appraisal profiles (relative to the “no distortion” baseline) for all distortion classes.
@@ 4.3 Discussion
Figure 6: Appraisal profiles (relative to the baseline) for selected distortion classes.
@@ 5 Cognitive Reframing and Emotional Regulation
Figure 7: The shift in appraisal profiles after cognitive reframing. Positive values mean that median appraisal value increase after the reframing, while negative values mean indicate the decrease.
@@ 7 Conclusion
Using automatic textual analysis to make inferences about people’s mental states has ethical implications. In this work, we use an existing publicly available dataset whose authors reported considering the ethical aspects of their data collection and annotation and also sought approval for their procedures from their institution’s review board (see (Sharma et al., 2023b) Section 4.3). Also, in our study, we are not developing systems for making predictions about individual persons, but rather studying the general patterns over groups.

### v2

@@ 1 Introduction
Emotions are expressed through various modalities, including tone of voice, facial expressions, gestures, and language, particularly in written text. This multifaceted expression of emotions has attracted the interest from NLP and computational researchers in recent years Wang et al. (2022); Plaza-del Arco et al. (2024). While discrete emotional states such as anger, joy, and fear are deemed universal and thus form the basis for automated emotion recognition research, a smaller number of studies have explored dimensional models, representing discrete emotions in continuous spaces Plaza-del Arco et al. (2024). Appraisal theories define emotions as responses that arise from an individual’s evaluation of and event’s significance to their personal goals and well-being, emphasizing that the quality and intensity of emotional responses depend on appraisals, which are the subjective interpretations of the situation Moors et al. (2013). In contrast to discrete emotional categories, appraisal theories maps an individual’s emotional state to a continuous space with each dimension representing an appraisal dimension. This not only provides a more detailed understanding of a person’s state, but also allows comparison between emotions.
Negative thoughts are a natural part of human experience; however, they can have a more profound impact on individuals with mental disorders, often becoming entrenched, automatic, and emotionally triggering. Cognitive distortions refer to irrationally exaggerated negative assessments of oneself or situations Beck (1963) and they are linked to the states of depression (Joormann and Stanton, 2016) and anxiety (Yazici-Çelebi and Kaya, 2022). Moreover, cognitive distortions have been found to correlate with the use of non-adaptive emotion regulation strategies (Deperrois, 2022). Cognitive restructuring, also known as cognitive reframing, is a therapeutic intervention designed to encourage a more positive outlook towards situations by addressing these negative thought patterns Clark (2013). This technique involves replacing negative thoughts with more neutral or hopeful “reframed thoughts”, which provide a softer alternative perspective on the situation.
@@
Our aim in this work is to bring together emotion appraisals and cognitive distortions, to explore the link between these two different but related psychological constructs. We believe that such a relation (if it exists) could be exploited to define more robust systems for automated emotion regulation. For instance, understanding of such relations could help to devise more deliberate and personalized ways for encouraging cognitive reframing and/or emotional reappraisal. We begin by training appraisal prediction models to perform automated appraisal annotations on a dataset annotated with cognitive distortions enabling a combined analysis of both constructs. We analyze the distribution of appraisal values for each distortion-appraisal pair individually, and find statistically significant relations between cognitive distortions and appraisal dimensions, suggesting that different distortion patterns may exhibit distinct appraisal profiles. Finally, when comparing the appraisal profiles between original and reframed texts, we observed a considerable positive shift in several appraisal dimensions, further demonstrating the link between the two constructs and supporting the need for combined study of the two areas.
@@ 2.4 Results
Our model performs on par with the results from Troiano et al. (2023). We use the root-mean-squared error (RMSE) as the evaluation metric and report macro-RMSE averaged over all appraisal dimensions on the test set of 1.36 compared to 1.40 reported in the original paper. Furthermore, Figure 2 compares the performance of the two models for each appraisal dimension, with our model out-performing Troiano et al. (2023) for 13 out of 21 appraisal dimensions. Figure 2 also reports the RMSE of just predicting the median ratings for individual appraisal dimensions, with both trained models performing better than the baseline median predictor in most dimensions. The average RMSE of the median predictor was 1.55, which is clearly worse than the trained models. Thus, we deem our trained model good enough to be used for automated appraisal annotation in the remainder of the paper.
Figure 2: Appraisal ratings prediction accuracy in RMSE for our model, the results reported by Troiano et al. (2023), and those calculated against median predictions for each appraisal dimension as baseline.
@@ 3.1 Thinking trap Dataset
Table 1: Statistics of the cognitive distortion labels of our version of the Thinking Trap dataset.
@@ 3.3 Negative groups
One major consideration is the definition of the negative group within this analysis. The relationship between cognitive distortions and emotional appraisal dimensions, as inferred from the Mann-Whitney test, is strongly influenced by how the positive and negative groups are defined. Although the definition of the positive group p is fixed, the negative group can have different meanings. Therefore, we consider the following three different definitions of the negative group n:
No distortion: in this case, the negative group only contains entries without any distortion, i.e., those belonging to the “no distortion” class. This group represents the appraisal profile of texts without cognitive distortions and acts as a global baseline against which we can compare individual distortion profiles.
Exclusive: here, the negative group contains entries that do not belong to the given distortion class (defining the positive group) but belong to other cognitive distortion classes (excluding “no distortion”). By utilizing this negative group, we can identify differences in appraisal values between various distortion classes. This approach enables us to analyze the appraisal profile of a specific distortion in relation to other distortions, rather than comparing it to a global baseline.
All others: in this configuration we combine both settings by including in the negative group all entries that do not belong to the specified distortion class (defining the positive group).
The remainder of the paper mostly focuses on the first two categories of negative groups since behavior of all others and exclusive categories was found to be identical in all our experiments. This can be attributed to the fact that roughly 90% of the data points express a distortion, thereby dominating the behavior of the “no distortion” class.
@@ 3.4 Results
Figure 3 plots the discretized significance values for two definitions of the negative groups considered in our study. [^6] Please refer to Appendix A for the remaining plots. In these plots, a cell colored white represents a statistically significant difference in appraisal values between the positive and negative groups, while a black cell indicates no statistically significant difference.
@@ 4 Distortions and Corresponding Appraisal Profiles
Figure 5: Appraisal profiles (relative to the “no distortion” baseline) for all distortion classes. The x-axis represents different appraisal dimensions considered while the y-axis plots the difference in appraisal scores between individual distortions and the baseline ($score(distortion)-score(baseline)$).
@@ 4.3 Discussion
Figure 6: Appraisal profiles (relative to the baseline) for selected distortion classes. The x-axis represents different appraisal dimensions considered while the y-axis plots the difference in appraisal scores between individual distortions and the baseline ($score(distortion)-score(baseline)$).
@@ 5 Cognitive Reframing and Emotional Regulation
Figure 7: The shift in appraisal profiles after cognitive reframing. Positive values mean that median appraisal value increased after the reframing, while negative values mean indicate a decrease.
@@ 7 Conclusion
Ethical Considerations
Despite the growing use of NLP (and AI in general) for analyzing the mental and emotional state of individuals based on a variety of input data sources, some ethical considerations need to be taken into account within the research process.
One major area of consideration is the data collection process for such studies. In this work, we use an existing publicly available dataset whose authors reported considering the ethical aspects of their data collection and annotation and also sought approval for their procedures from their institution’s review board (see (Sharma et al., 2023b) Section 4.3).
This research direction also comes with significant ethical considerations pertaining to the use of such models. Despite the growing interest in automated systems for mental health analysis and monitoring, improper use of these systems can lead to issues like labeling and stigma. Within our study, we are not developing systems for making predictions about an individual’s mental health, but rather studying the general patterns over groups.